| <p>On 23/02/2018:</p>
<p>At the end of the February 2018 shifts the detectors bias settings were:</p>
<p>S2 bias = -130.3V, leak I = 2.42uA</p>
<p>Pin diode bias = -70.1V, leak I = 0.00uA.</p>
<p>Both detectors were switched off, the preamp power was also switched off.</p>
<p>At 18:22 the chamber pressures read:</p>
<p>TPU1 = 2.3e-6 mBar</p>
<p>TPU2 = 6.9e-7 mBar</p>
<p>TPD1 = 1.6e-6 mBar</p>
<p>TPD2 = 1.7e-6 mBar</p>
<p>CHUP = 5.7e-6 mBar</p>
<p>CHDW = 1.5e-5 mBar</p>
<p> </p>
<p>We closed the manual gate valve and T1600 Backing Valves. We switched off the turbos and roughing pumps Ecodry 1 and 2.</p>
<p>The nim bin for the quad bias and other nim modules was switched off. The Lecroy 1434A CAMAC Crate was switched off. The RAL shaping amp crate was switched off.</p>
<p>We put away the S2 and pin diode detectors.</p>
<p>The alpha source was removed from the target ladder.</p>
<p>The target ladder was remounted in the following configuration:</p>
<table border="1" cellpadding="1" cellspacing="1" style="width:500px">
<tbody>
<tr>
<td>Position</td>
<td>Target</td>
</tr>
<tr>
<td>1</td>
<td>Blank 3mm</td>
</tr>
<tr>
<td>2</td>
<td>
<p>No Target</p>
</td>
</tr>
<tr>
<td>3</td>
<td>CH2 VII</td>
</tr>
<tr>
<td>4</td>
<td>CH2 III</td>
</tr>
<tr>
<td>5</td>
<td>CH2 IV</td>
</tr>
<tr>
<td>6</td>
<td>No Target</td>
</tr>
<tr>
<td>7</td>
<td>No Target</td>
</tr>
</tbody>
</table>
<p>We performed a general tidy-up around the chamber and workspaces we used during the past 9 days.</p>
<p>I took the PB-5 Pulse Generator and three 64 Way - 2X32 Way ribbon cables (for minor modification) back to Edinburgh.</p>
<p>I also took the mylar we discovered in the beamline on Tuesday 20th as a keepsake/ reminder anything can happen</p>
<p>At ~20:00 we declared victory for the week.</p> |